OpenStudy (anonymous):

A rotor is run at the following speeds for intervals of 1 minute: |Minute: | 1 | 2 | 3 | 4 | |Speed: | 1 | 2 | 4 | 8 | How many minutes will it take to reach a speed of 64 meters per second? a.) 6 b.) 7 c.) 8 d.) 9 Thank you oh so very much. <3 Xoxoxo.

OpenStudy (welshfella):

Hint; the speed is doubling every minute

OpenStudy (anonymous):

finish the chart. 1 - 1 2 - 2 3 - 4 4- 8 5 - 10 6 - 12 7 - 14 8 - 16 9 - 18

OpenStudy (anonymous):

Lets make it simple. C

OpenStudy (anonymous):

10 - 20 11 - 22 12 - 24 14 - 26 15 - 28 16 - 30 18 - 32

OpenStudy (anonymous):

19 - 34 20 - 36 21 - 38 22 - 40 ugh.. this is tedious. :( lol

OpenStudy (anonymous):

Your answer is C lol

OpenStudy (anonymous):

by the time u reach 64 you will be at your 8th interval. so u could just multiply :/

OpenStudy (anonymous):

Lol, sorry. I know I look like an idiot. For some reason math just doesn't click... like at all for me. Haha.. Thank you though. ^_^

OpenStudy (welshfella):

after 5 seconds the speed is 2*8 = 16 after 6 s it will be 2*16 = 32 after 7 s what would the speed be?

OpenStudy (anonymous):

48 >_>

OpenStudy (welshfella):

the answer is not 8

OpenStudy (welshfella):

note the common factor is 2 each term is obtained from previous by multiplying by 2

OpenStudy (anonymous):

8 is the closes man. or girl... idk

OpenStudy (welshfella):

man

OpenStudy (anonymous):

okay cool

OpenStudy (anonymous):

Ooooh! So it's 2 * 32!

OpenStudy (welshfella):

this sequence is called a geometric sequence with common ratio 2

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

Yaaay so it's 7!

OpenStudy (welshfella):

7th term = 2*32

OpenStudy (welshfella):

yea

OpenStudy (anonymous):

Gah! Thank you for your patience with me. It's greatly appreciated :)

OpenStudy (welshfella):

yw